# Assembly Theory Reduced to Shannon entropy and Rendered Redundant by Naive Statistical Algorithms

### v7

@@ 2.6 Block-decomposing the Assembly Index
It is straightforward to demonstrate that the Ai can never perform better than BDM, since the latter counts exact copies in data as its most basic test for non-randomness, which means that Ai is at BDM’s baseline. However, unlike Ai, BDM will penalise for trivial repetitions that are also found in simple stochastic processes, such as rock formations or fractal-like crystals [7], unlike Ai.
As discussed in Section 4, we demonstrated in [4] that the $\mathbf{LZ}$ parsing of any object according to the assembly index (Ai) calculation method is given by the longest (rooted) assembly paths that belong to the minimum rooted assembly subspace from which the assembly index of the object is calculated.
@@
Then, it is known that the best statistical compression rate by any method is lower bounded by entropy-based compression. Statistical compression methods such as Huffman, LZW or LZ78 are universal in that can they achieve such a Shannon entropy ($\mathbf{H}$) rate limit [15].
@@
where $K_{Ai}\left(\cdot\right)$ is the size in bits of the compressed form according to the assembly index calculation method.
In the general case, one can also directly demonstrate that the assembly index calculation method can never outperform BDM for any generative process (as opposed to simple stochastic sources), whether one has a fully deterministic source, stochastic source, or a mixed process. We have demonstrated in [4] that any a priori chosen encoding scheme $V\left({\Gamma^{*}}_{y}\right)\right>$ of the set $V\left({\Gamma^{*}}_{y}\right)$—an encoding which is agnostic not only vis-a-vis the choice of the final object but also the (sub)space structural representation—of vertices of the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ of the (final) object $y$ losslessly compresses/encodes the object $y$. Thus, there is a partition $i$ sufficiently close to $i_{0}$, and there is an encoding-decoding scheme $m$—under computational resources of the same order as those expended to calculate the assembly index—such that
@@
where |$V\left({\Gamma^{*}}_{y}\right)\right>\right|$ | is the size in bits of the arbitrarily chosen encoded form of the set $V\left({\Gamma^{*}}_{y}\right)$. Thus, Ai is a compression method subsumed within BDM (see also Equation (5)).
@@
$-\log\left(\sum\limits_{\mathbf{U}\left(p\right)=x}\frac{1}{2^{\left|p\right|}}\right)$ $=$ (3)
$\mathbf{K}\left(x\right)\pm\mathbf{O}(1)$$\leq$
@@
where $\sum\limits_{\mathbf{U}\left(p\right)=x}2^{-\left|p\right|}$ is the universal a priori probability—or, equivalently, the algorithmic probability (AP)—of the event $x$.
Another core concept in AIT is algorithmic randomness, of which statistical randomness–that is, randomness that can be verified by a statistical test, such as Shannon entropy-based indexes– is a particular (proper) case. The halting probability $\Omega\in\left[0,1\right]\subset\mathbb{R}$ (also known as Chaitin’s constant or Omega number), defined by
$\Omega=\sum\limits_{\begin{subarray}{c}\exists y\left(\mathbf{U}(p)=y\right)\end{subarray}}\quad\frac{1}{2^{\left|p\right|}}\text{ ,}$
is known to be incompressible (or algorithmically random), i.e.,
$\mathbf{K}\left(\Omega\upharpoonright_{n}\right)\geq n-\mathbf{O}(1)$
holds for a sufficiently large $n$. Not only is this number the canonical example of algorithmic randomness, passing every formal-theoretic or computable statistical test that one may conceive, but it also subsumes all the information (in its most compressed form) necessary to solve the halting problem. In this manner, both the algorithmic probability or the algorithmic complexity of any object can always be computed if the first bits of $\Omega$ (denoted by $\Omega\upharpoonright_{n}$) are given as input. Formally, there is a fixed program $p$ of a universal machine $\mathbf{U}$ such that for any $x$,
$\mathbf{U}\left(x,\Omega\upharpoonright_{\left|x\right|+\mathbf{O}\left(\log\left(\left|x\right|\right)\right)},p\right)=\mathbf{K}\left(x\right)\text{ .}$ (4)
Contrary to multiple false claims by the authors of Assembly Theory and their limited examples allegedly proving divergence from popular statistical compression [1], the present article empirically investigated the asymptotic behaviour of BDM, Ai, and LZW for those cases confirming Equations 1 and 2.
@@
Ai not only offers no advantage over BDM (or even $\mathbf{H}$ or LZW) but also no computational complexity advantage, as its calculation, according to the authors themselves, is intractable (a fact they present as an advantage) [1]. For this reason they have also had to come up with different heuristics for every paper they have published before [5, 23, 3], even when their algorithm only counts naive statistical redundancies (exact copies).
A naive BDM implementation capable of only counting copies (with no CTM advantage) would run in constant to linear time depending on implementations (just as with $\mathbf{LZ}$ algorithms) and would then approximate $\mathbf{H}$ just as Ai does [7, 4]. In contrast, any run of CTM would give BDM an advantage Ai would never have at finding causal redundancy in an object, beyond the triviality of finding simple exact copies [8, 9].
@@
For the sake of denoting a hierarchical relationship between complexity methods and complexity indexes, let $A\preceq_{cp}B$ denote the (partial) order of a method $B$ subsuming another complexity method $A$ (and $\equiv_{cp}$ when both $A\preceq_{cp}B$ and $B\preceq_{cp}A$ hold). Thus, from [4] and Equations (15), (1), (2), (3), and (4), one has it that the following relationship
$\textnormal{Ai}\equiv_{cp}\textnormal{LZ}\preceq_{cp}\mathbf{H}\prec_{cp}\textnormal{BDM}\preceq_{cp}\textnormal{CTM}\prec_{cp}\textnormal{AP}\equiv_{cp}\mathbf{K}\preceq_{cp}\Omega$ (5)
holds.
@@
Additionally, BDM is computable, just as Ai (or, equivalently, LZ) is, but Ai is a weakest loosest bound unable to capture any non-trivial pattern or causal content and it cannot explain selection or evolution based on counting copies and circular arguments [4] any better than $\mathbf{LZ}$ and Shannon entropy can [2].
@@
Equation 5 summarises some of the most important historical and methodological concepts in the history of complexity theory, creating a hierarchical structure of complexity measures into which Ai was never properly inserted, the senior authors even suggesting that Ai figured at the rightmost end as a generalisation of all other complexity measures [20]. Not only is that wrong, with Ai arriving probably 50 years late, but Ai implements the weakest form of test that any other measure of complexity would already take as its baseline. In fact, the same copy number argument has been used in the same context for decades, even in the reconstruction of evolutionary trees [31, 32] and in the addressing of some of the greatest challenges in biology, such as nucleosome positioning [33] in connection with GC content (counting nucleotides G and C). We made this case in a set diagram in [4], here formalised in Equation 5. Every measure formulated to date can be inserted somewhere between these lower and upper bounds.

### v10

@@ 2.2.1 Analysis on Synthetic Strings
2.2.2 Analysis on Molecular Data
To explain how almost any naive algorithm can reproduce the results reported by the authors of Assembly Theory in [1, 2], we took the data from [1] and reproduced their analysis in Fig. 4 by exploiting the linear relationships found between LZW, the length of the InChI strings for molecules, and MS2 data, respectively.
We say that the length of the InChI string is the number of characters existing after the initial prefix “InChI=1S/”. We also plot in Fig. 4 the relation to MA (this is, Ai), as presented by Marshall et al. [1] in their Fig. 3. In such a paper, they call their Ai, ‘MA’, standing for Molecular Assembly; they are for all purposes the same metric based on counting molecular copies. Ai is a generalisation of MA that the authors have applied to objects other than molecules while MA was applied exclusively to molecules. From now on, we will therefore use MA and Ai interchangeably, but will use MA in the context of their first paper applied to molecules from which they claim to be not only connected but able to fully explain selection and evolution.
Figure 4: Quantile regressions considering 5%, 50 % and 95% quantiles and Pearson correlation (corr) between MA (top), Length of InChI string (middle) and LZW of InChI string (bottom) and MS2 using the data from [1]. The shaded area represents a 90 % quantile interval. For any of the metrics, linear relationships with similar prediction capabilities are observed. This directly undercuts the AT authors’ claims that one of the main advantages of MA (or, equivalently, Ai) is that it could be experimentally measured, since all the other metrics (LZW and the Length of the InChI codes) can also be estimated from experimental MS2 data directly, as we have done here. In contradiction to what the authors claim in [4], the authors of AT have failed at controlling for molecular length, which drives their index. The authors of AT have thus far failed to show a single example in which LZ or Shannon entropy cannot reproduce any results produced by Ai or AT, meaning that Ai and AT are simply equivalent to applying LZ or Shannon entropy with many extra steps.
One might ask why plotting the relation between MA and the length of InchI strings is of interest. This is a consequence of the asymptotic behaviour observed for compression rates in LZ schemes, which tend to a constant for large strings. Since Ai belongs to the LZ family [17], Ai should be linearly related to the length of the base string for which it is calculated. Thus, it would be natural to observe some type of linear relation, which was verified in Fig. 5.
Figure 5: Linear relation between MA and the Length of the InChI strings of molecules using the data from the original AT paper [1]. This shows that just by taking the length of SMILES notation, one can reproduce the results of AT and their assembly index. If AT has any advantage over other representations and other compression algorithms, it should produce different if not better results, but it does not. This was conducted on the same set of molecular compounds used by the authors in their paper [1] as the basic control experiment they did not perform, and was reported before to be able to separate organic from non-organic molecules [22].
As shown in 5, the authors of AT did not control for molecular length, which would have deleted any effect of Ai as it does not pick up any other signal.
Figure 6: Values of several complexity measures for Small Molecules and Peptides, using the data from [1]. This is the calibration experiment performed in [1] before moving to classify organic vs nonorganic compounds. While in [22] we normalised by molecular length and compared the efficiency of the organic vs inorganic separation on over 15,000 compounds reported (the full ChemicalData[] database from the Wolfram language), not a small selection [1] of only about 130 compounds, with no comparisons to any other complexity measure that would have disclosed similarity or superiority at separating the classes. MA is one of the many versions of Ai across the authors’ papers for application to molecules, implementing their ‘copy-number’ hypothesis.
From our numerical experiments it is clear that LZW and Ai produce results that cannot be told apart. In Fig. 2, a clear linear asymptotic trend is observed between the LZW and Ai values. This suggests that one can easily derive one from the other except for small local variations that vanish rapidly as a function of length.
Fig. 3 shows that the mean value of (log LZW/log Ai) tends to a constant and that the standard deviation of (log LZW/log Ai) tends to 0 for large strings. This indicates that the random variable (log LZW/log Ai) tends to a single-bar histogram, indicating that this ratio tends to a constant without variability. This is an empirical demonstration that the two values converge in alignment with the mathematical proof of equivalence provided in [17].
Fig. 4 shows that the claimed link that Ai has with the experimental results is also observed for LZW and even for the length of the InChI strings of the molecules analysed. Thus, if we follow the same rationale as Marshall et al. [1], it could be said that LZW of the InChI strings, and even on their lengths only, can be experimentally measured by molecules’ mass spectra. This lacks, of course, a physical/chemical explanation, which is never provided by AT, beyond circular arguments [17].
Also, taking into account the results of the experiments in Fig. 2, one could investigate how the Ai values are related to the length of the InChI strings. Fig. 5 shows a remarkably strong linear relationship, with an $R^{2}$ of more than 0.9 and a Pearson correlation of 0.95. This indicates that Ai is a proxy for the length of the InChI strings, revealing that these strings already present a great share of the “concatenation” information for the molecules.
Regarding the threshold suggested by [1], that according to the authors separates molecules produced by nonliving systems from molecules produced by living ones, it is claimed that nonliving systems do not produce molecules with MAs (Ai) higher than 15. This has already been found to be erroneous by two other independent groups  [23, 24]. However, in their defence [4] in this regard, the authors suggest that one has to filter out those examples that prove the Ai hypothesis wrong because they are too artificial, while at the same time, the same authors claim that Ai is completely agnostic and capable of detecting any form of life in the universe [1, 2].
This claim follows from [1, Fig. 4], where it is asserted that only molecules produced by living organisms (and some “Blinded” samples from NASA/other sources that were not made fully available in their paper’s dataset, bypassing Nature’s rules and review procedures) have MAs (Ai) higher than 15. By using their own data, we reproduced their [1, Fig. 4] in our Fig. 7, taking the results for Ai at face value as they were not reproducible, given that despite their claims to the contrary, they did not include all the data. The threshold has been already challenged and found to be in error twice by other independent groups [23, 24], but here we show clearly that their assembly index (Ai) is not picking up any special signal from life.
The first fact worth noting is that Marshall et al. [1] did not use any of the quantile regressions they indicated in their [1, Fig. 3] to estimate MA (Ai) values from MS2 peaks (i.e., they did not do any of the 5%, 50 % and 95% quantiles, which are also reproduced in Fig. 4). By experimenting with the quantile value, it was possible to determine that they used a 14% quantile regression on the data to build their linear relationship. Then, using this linear relationship and the MS2 peak measurement data, Fig. 7 was obtained. This choice was not explored in depth in their paper, indicating a serious methodological gap.
If we follow the same rationale as in [1], we could say that molecules produced by nonliving systems could not have estimated InChI string lengths greater than 100 characters. This illustrates why their claim is neither methodologically robust nor correct. Thus, the analysis of Figs. 4 and 7 shows that the organic vs inorganic separation claimed in [1] is driven by molecular length and not any other feature captured by their (molecular) assembly index or their copy number, beyond triviality.
We reported all this years before AT in [22], using a more comprehensive experiment involving over 15 000 compounds, and with proper control experiments by comparing to several measures and several data representations, and without offering the results as proof of unification of physics and biology [2]. The rationale behind Assembly Theory is equivalent to the most common practice of finding any variable linearly correlated to MS2 and then using an arbitrary threshold to separate the molecules, which is without any physical/chemical foundation and does not withstand basic scrutiny.
The authors of AT have unwittingly fallen into the classic correlation as a causation trap, where one finds a correlated variable and claims a causal effect.
Despite the decades of research showing the application of information theory and compression methods to applied sciences, Assembly Theory’s proponents keep offering up the fallacy—as in the abstract of [18]—that compression methods are not efficiently applicable, or in general that algorithmic complexity-based measures cannot be applied to empirical data or in experimental applications. In [18], the Molecular Assembly index (based on the assembly index) was shown to correlate (with a Pearson’s of 0.88 for their best result) to an ensemble of experimentally measured data. If this correlation “demonstrates” that Assembly Theory is an experimental measure, then it follows that the strong correlation using other metrics we have shown would also “demonstrate” that one can measure ‘assembly’ by using LZ or Shannon entropy.
Since the relation between MS2 peaks and MA (Ai) is linear (the same for InChI lengths), what was actually proposed in one of the papers in question [1] is that molecules produced by living and nonliving systems can simply be distinguished by the number of their MS2 peaks (which were measured and not estimated, the authors contend).
2.3 AT Unsuitability to Quantify Selection or Evolution
One of the recurring difficulties in engaging with Assembly Theory (AT) is the way its proponents shift domains when their claims are challenged. At the molecular level, the authors of AT insist that the assembly index (Ai) captures the causal essence of biological selection. Yet when it is shown that Ai fails—where Shannon entropy and naive compression already replicate or outperform its results—they retreat to the claim that Ai is broader, extending beyond molecules and even beyond biological evolution. This rhetorical oscillation highlights the absence of a consistent explanatory framework. A measure that cannot withstand scrutiny in its supposed “home domain” of molecular complexity cannot be rehabilitated by vague appeals to universality.
Here we offer some arguments showing how AT with its Ai would fail in characterising simple processes of biological evolution and selection. Let $o$ be any object (molecular or not) and $x=\mathrm{enc}(o)\in\Sigma^{*}$ its symbolic encoding. Let $Ai(x)\in\mathbb{N}$ denote the assembly index of $x$, defined as the minimal number of exact-copy concatenations required to generate it. Let $g$ be a generative description of $o$, and let the realised phenotype be $\phi(g,o)$. In an environment $e$, fitness is given by $w(\phi(g,o),e)$. For a trait $\tau$ with indicator $\mathbf{1}_{\tau}\in\{0,1\}$, the standard selection coefficient is
$s_{\tau}(e)=\mathbb{E}[w(\phi(g,o),e)\mid\mathbf{1}_{\tau}=1]-\mathbb{E}[w(\phi(g,o),e)\mid\mathbf{1}_{\tau}=0].$ (2)
Trait loss (regressive evolution) occurs when the net benefit $b_{\tau}(e)$ of $\tau$ falls below its maintenance cost $c_{\tau}$, so that $b_{\tau}(e)-c_{\tau}<0$.
  1. 1.
    Environment dependence. For fixed $x$ (hereinafter fixed $Ai(x)$), choose environments $e_{1},e_{2}$ such that $s_{\tau}(e_{1})>0$ while $s_{\tau}(e_{2})<0$. Since $Ai(x)$ does not vary with the environment, no function of $Ai(x)$ can uniformly recover the sign of $s_{\tau}(e)$.
  2. 2.
    Trait-loss edits invisible to Ai. Let $T$ be a causal transformation (deletion, reversion, inversion, simplification) yielding $o^{\prime}=T(o)$ with encoding $x^{\prime}=\mathrm{enc}(o^{\prime})$. Such edits can eliminate costs ($c_{\tau}\to 0$) without introducing exact-copy patterns. Hence $Ai(x^{\prime})$ may change arbitrarily, while the true selective dynamics is driven by causal edits invisible to Ai.
  3. 3.
    Impossibility. If there existed $F$ with $s_{\tau}(e)=F(Ai(x))$ for all $x,e$, then by (1) the same $Ai(x)$ would imply contradictory signs of $s_{\tau}(e)$. Therefore, no such $F$ exists. Selection coefficients cannot be functions of Ai (or any statistic of isolated encodings) alone.
Whether $o$ is molecular or not, selection—and de-evolution in particular—depends on $(o,e)$, not on $o$ in isolation. Ai, as a function of isolated encodings, cannot capture the environment-dependent and causally structured nature of evolutionary dynamics. Thus, Ai cannot be a good metric for selection or evolution if it cannot explain even its most basic operations.
The Cavefish and the Limits of Assembly Theory
The Mexican blind cavefish (Astyanax mexicanus) provides a concrete biological case that exposes the conceptual and empirical limits of Assembly Theory (AT) and its Assembly Index (Ai). This species has evolved multiple independent cave-dwelling populations derived from surface ancestors, exhibiting striking regressive and adaptive traits such as loss of pigmentation, eye degeneration, expanded taste buds, and rewired sensory systems. These transformations are driven by well-understood evolutionary mechanisms including natural selection, relaxed selection, and environmental constraint—all processes that are inherently causal, reversible, and context-dependent.
Such systems are directly relevant to AT because their authors explicitly claim that their framework ‘reconstructs the evolutionary tree of life’, ‘captures causal molecular relationships’, and ‘maps evolutionary trajectories’ from only molecular data. If AT genuinely encodes causal history and evolutionary change, it should at least accommodate these canonical cases where adaptation and trait loss are understood in molecular and developmental terms. The cavefish therefore serves as a clear empirical test for AT’s central claim.
Table 4: Evolutionary mechanisms in cavefish and the limits of Ai/AT, contrasted with the broader capabilities of BDM/CTM.
Evolutionary mechanism Biological example Why Ai/AT fails How BDM/CTM may capture the traits
Environment-dependent selection Eyes are adaptive in light-rich rivers, but maladaptive in dark caves Ai assigns a fixed value to molecular encodings, blind to environmental context BDM/CTM incorporate environment via algorithmic probability and reweighting of generative models
Regressive evolution (trait loss) Eye reduction and disappearance saves energy in darkness Trait loss involves deletions and simplifications, not repetitive motifs (even at the molecular level) BDM accounts for deletions, reversions, and inversions; CTM explores alternative generative programs reflecting simplification
Selection not reflected at molecular level Enhanced taste buds, mechanosensory cells, and altered foraging behaviours Adaptive traits arise through regulatory/neural changes, invisible to Ai’s copy-counting BDM captures causal, non-statistical regulatory patterns; CTM identifies minimal generative rules beyond copy-number
The cavefish example demonstrates that evolution is not a monotonic accumulation of structural novelty or statistical regularity. It involves gains, losses, and shifts in function that depend on the ecological context and developmental regulation. Ai reduces all these processes to an invariant count of combinatorial steps, blind to the mechanisms by which traits appear or disappear. Even at the molecular level, most adaptive changes in Astyanax occur through altered gene expression and sensory integration rather than through the creation of new metabolites. Consequently, any approach restricted to molecular assembly complexity would miss precisely the causal signals it claims to capture.
Moreover, when AT is applied to actual molecular datasets, its results are indistinguishable from those produced by simple surrogates such as Shannon entropy, GC content, or compression-based similarity. These established information-theoretic measures already reconstruct molecular evolutionary trees with high accuracy and known statistical properties. If AT cannot outperform these benchmarks, its claim to causal or evolutionary insight collapses into rebranded statistics.
The cavefish illustrates the broader conceptual failure of AT’s central premise. The theory conflates structural description with evolutionary causation, ignoring the roles of environment, regulation, and developmental plasticity. If Ai cannot account for the causal mechanisms that drive adaptation in such a textbook case of evolution, then its promise to explain selection and evolution from molecular data alone cannot be sustained. The example therefore stands not as an isolated counterpoint but as representative of the many causal, non-trivial features of evolution that AT is structurally incapable of detecting.
Ai reduces selection and evolution to trivial statistics, missing the causal mechanisms at play that do not lead to statistical clues not reflected in the static outcome of the object at a given time, despite the many suggestions that AT and Ai capture the dynamic causal history of the object.
This example undermines AT’s pretension of generality or relevance in the very core area in which they have introduced it. Even at the molecular level, Ai performs no better than simple surrogates, as we have shown, the raw length of molecular encodings such as InChI strings and the application of Shannon entropy to their own data reproduces their results. If Ai fails in its core claim of detecting selection from molecular data, extending it before and beyond molecules, but it cannot reproduce any new result, it cannot be presented as a new theory but rather as a pedagogical exercise.
In contrast, measures grounded in algorithmic information theory, such as the Block Decomposition Method (BDM), introduced before AT, subsume trivial copycounting while are also capable of capturing other causal operations ubiquitous in biology such as deletions, reversions, inversions, and other structured mechanistic transformations.
@@ 2.4 Block-decomposing the Assembly Index
As detailed in Section 4, the Block Decomposition Method (BDM) offers a more robust framework for complexity estimation than the assembly index (Ai). By its construction, Ai can never perform better than BDM because Ai’s core mechanism—counting exact copies—is merely the baseline test for nonrandomness that BDM incorporates.
As discussed in Section 4, we demonstrated in [17] that the LZ parsing of any object according to the assembly index (Ai) calculation method is given by the longest (rooted) assembly paths that belong to the minimum rooted assembly subspace from which the assembly index of the object is calculated.
@@
Then, it is known that the best statistical compression rate by any method is lower bounded by entropy-based compression. Statistical compression methods such as Huffman, LZW or LZ78 are universal in that can they asymptotically all converge to Shannon entropy ($\mathbf{H}$) rate limit [25].
@@
where $K_{Ai}(x)$ is the size in bits of the compressed form according to the assembly index calculation method.
In the general case, one can also directly demonstrate that the assembly index calculation method can never outperform BDM for any generative process (as opposed to simple stochastic sources), whether one has a fully deterministic source, stochastic source, or a mixed process. We have demonstrated in [17] that any a priori chosen encoding scheme $\left<V\left({\Gamma^{*}}_{y}\right)\right$> of the set $V\left({\Gamma^{*}}_{y}\right)$—an encoding which is agnostic not only vis-a-vis the choice of the final object, but also the (sub)space structural representation—of vertices of the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ of the (final) object $y$ losslessly compresses/encodes the object $y$. Thus, there is a partition $i$ sufficiently close to $i_{0}$, and there is an encoding-decoding scheme mm—under computational resources of the same order as those expended to calculate the assembly index—such that
@@
[…]
@@
$-\log\left(\sum\limits_{\mathbf{U}(p)=x}\frac{1}{2^{|p|}}\right)$ $=$ (5)
$\mathbf{K}(x)\pm\mathbf{O}(1)$ $\leq$
@@
where $\sum_{\mathbf{U}(p)=x}2^{-|p|}$ is the universal a priori probability—or, equivalently, the algorithmic probability (AP)—of the event $x$.
Contrary to multiple claims by the authors of Assembly Theory and their examples [4], we have empirically shown the asymptotic behaviour of Ai to other metrics, including LZW.
@@
Ai offers no advantage over BDM, Shannon entropy ($\mathbf{H}$) or LZW, but also no time complexity advantage, as its calculation, according to the authors themselves, is intractable (a fact they present as an advantage) [4]. For this reason, they have also had to come up with different heuristics in different papers [1, 26, 2].
@@
The following equation provides a hierarchical relationship between complexity methods to locate the place of each metric. Let $A\preceq_{cp}B$ denote the (partial) order of a method $B$ subsuming another complexity method $A$ (and $\equiv_{cp}$ when both $A\preceq_{cp}B$ and $B\preceq_{cp}A$ hold). Thus, from [17] and Equations (16), (3), (4), and (5), one has it that the following relationship
$\textnormal{Ai}\equiv_{cp}\textnormal{LZ}\preceq_{cp}\mathbf{H}\prec_{cp}\textnormal{BDM}\preceq_{cp}\textnormal{CTM}\prec_{cp}\textnormal{AP}\equiv_{cp}\mathbf{K}$ (6)
holds. Where Ai and $\mathbf{K}$ are at the two extremes of the complexity spectrum [27] and Ai on the trivial side, weaker or equal to popular statistical compression and Shannon entropy as only counting repetitions, while computable approximations related to stronger measures have been proven to better capture causality [28] related to selection and evolution [29].
@@
In addition, BDM is computable given the relaxation of the condition of not having to find or prove program length minimality (or, equivalently, all the generating programs). Ai is the weakest metric, unable to capture any nontrivial pattern, let alone nontrivial causal content; therefore, it cannot explain selection or evolution despite its circular arguments [17], especially not any better than LZ and Shannon entropy can [5].
@@
Equation 6 summarises some of the most important historical and methodological concepts in the history of complexity theory, creating a hierarchical structure of complexity measures into which Ai was never properly inserted, the senior authors even suggesting that Ai figured at the rightmost end as a generalisation of all other complexity measures [16]. Not only is that wrong, with Ai arriving probably 50 years late, but Ai implements the weakest form of test that any other measure of complexity would already take as its baseline. In fact, the same copy-number argument has been used in the same context for decades, even in the reconstruction of evolutionary trees [30, 31] and in the addressing of some of the greatest challenges in biology, such as nucleosome positioning [32] in connection with GC content (counting nucleotides G and C). We made this case in a set diagram in [17], here formalised in Equation 6. Every measure formulated to date can be inserted somewhere between these lower and upper bounds.